# A multiplicity formula for the Milnor number of smoothable curves

### v2

@@ 3. A flatness result
We apply Lemma 3.1 with $R=\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$, $I:=I_{X}$, $J:=I_{W}$, and $\mathfrak{m}$ the ideals of $X$, $W$ and 00 in $\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$, respectively. Note that $I_{X}\cap I_{W}=I_{Z}$ and the ideal of $S$ in $\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$ is $I_{X}+I_{W}$. Let $t$ be the uniformizing parameter of $(\mathcal{O}_{Y},0)$. Observe that the depth hypothesis of Lemma 3.1 are satisfied. Indeed, $Z$ is a complete intersection of dimension $2$, so $\mathrm{depth}(\mathcal{O}_{Z})=2$; $X$ is flat over $Y$, and so $t$ is a nonzero divisor of $\mathcal{O}_{X}$ and thus $\mathrm{depth}(\mathcal{O}_{X})\geq 1$. Also, because $I_{W}$ is obtained from the primary decomposition of $I_{Z}$ by removing some of its minimal primes, then $t$ avoids the minimal primes of $I_{W}$, and so $W$ is flat over $Y$. Thus, $\mathrm{depth}(\mathcal{O}_{W})\geq 1$. If $S$ is empty there is nothing to prove. So we can assume that either $S$ is set-theoretically equal to 00, or $S$ is of dimension one. The first case is impossible because by Lemma 3.1 $\mathfrak{m}\notin\mathrm{Ass}(\mathcal{O}_{S})$.
Suppose $S$ has a vertical component, i.e. $X_{0}$ and $W_{0}$ share an irreducible component. Identify $I_{X}$ and $I_{W}$ with their images in $\mathcal{O}_{Z}$. Suppose that as sets $X_{0}\subset W_{0}$. But $Z_{0}=X_{0}\cup W_{0}$, so $W_{0}=Z_{0}$. This means that for each $f\in I_{W}$ we have $f=tf_{1}$ with $f_{1}\in\mathcal{O}_{Z}$. Because $\mathcal{O}_{W}$ is flat over $\mathcal{O}_{Y}$, we have $f_{1}\in I_{W}$. So $I_{W}=tI_{W}$ in $\mathcal{O}_{Z}$. By Nakayama’s lemma $W=Z$, which is impossible.
Assume that $X_{0}$ is not contained in $W_{0}$ set-theoretically. Then there exists an irreducible component of $X_{0}$, which is not a component of $W_{0}$. Consider the union of all such components of $X_{0}$ that are not components of $W_{0}$ and denote it by $X_{0}^{\prime}$. Pick an element $r\in\mathcal{O}_{Z}$ such that its image $\overline{r}$ in $\mathcal{O}_{Z_{0}}:=\mathcal{O}_{Z}/t\mathcal{O}_{Z}$ is in the ideal of $X_{0}^{\prime}$ but not in the ideal of $X_{0}$. Then for each $f\in I_{W}$, there exists $f_{1}\in\mathcal{O}_{Z}$ such that $rf=tf_{1}$. Again, because $\mathcal{O}_{W}$ is flat over $\mathcal{O}_{Y}$, we have $f_{1}\in I_{W}$. Thus $rI_{W}\subset(t)I_{W}$. By the Cayley–Hamilton theorem $\det(A)I_{W}=0$ where $A=(\delta_{i,j}r-a_{i,j})$ is a square matrix of size $m\times m$ for some $m$, $\delta_{i,j}$ is Kronecker’s delta function and $a_{i,j}\in(t)$. But $I_{X}=((0)\colon_{\mathcal{O}_{Z}}I_{W})$. Thus $\det(A)\in I_{X}$. Denote by $\tilde{r}$ the image of $r$ in $\mathcal{O}_{X_{0}}$. Because $\det(A)$ is a monic polynomial in $r$ with coefficients in $(t)$, we have $\tilde{r}^{m}=0$ in $\mathcal{O}_{X_{0}}$. But this is impossible because we chose $r$ so that it’s not in the nilradical of $\mathcal{O}_{X_{0}}$. We reached a contradiction with the assumption that $S$ has a vertical component. Therefore, by taking sufficiently small representatives of the germs $(X,0)$ and $(Y,0)$ we can ensure that $(S,0)\rightarrow(Y,0)$ is finite (see [GrR84, Prp. 1), p. 63]).
Again, by Lemma 3.1, the origin cannot be an embedded component of $S$. Thus, $t$ is not a zero divisor of $\mathcal{O}_{S}$. So, $(S,0)\rightarrow(Y,0)$ is flat, and thus $y\rightarrow I(X_{y},W_{y})$ is constant for all $y$ close to 00. ∎
@@ 4. Proofs
Let $h\colon(X,0)\rightarrow(Y,0)$ be a one-parameter embedded flat deformation of $(X_{0},0)\subset(\mathbb{C}^{n},0)$ with $X_{y}$ smooth for small enough $y\neq 0$. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a complete intersection curve which contains $(X_{0},0)$ as a subvariety. Denote by $Z$ the induced deformation on $Z_{0}$. If $(X,0)$ is cut out from $(\mathbb{C}^{n}\times Y,0)$ by the vanishing of the analytic functions $f_{1},\ldots,f_{p}$, then the equations for $Z$ are given by $A(f_{1},\ldots,f_{p})^{T}$ for some $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$. Set $W:=\overline{Z\setminus X}$ and set $S:=X\times_{\mathbb{C}^{n}\times Y}W$. Observe that because $S$ is flat over $Y$ by Prp. 3.2, we have $W_{y}=\overline{Z_{y}\setminus X_{y}}$ for each $y\in Y$, i.e. the formation of $W$ specializes with passage to the fibers.
Denote by $J_{\mathrm{rel}}(X)$ the relative Jacobian module and denote by $J_{\mathrm{rel}}(Z):=AJ_{\mathrm{rel}}(X)$ the $\mathcal{O}_{X}$-module generated by the relative Jacobian matrix of $Z$. For each $y$ denote by $J(Z_{y})$ the image of $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{y}}^{n-1}$. It’s the $\mathcal{O}_{X_{y}}$-module generated by the columns of the Jacobian matrix of $Z_{y}$. We will repeatedly use the fact that the formation of Fitting ideals commutes with base change. In particular,
$\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))\otimes\mathcal{O}_{X_{y}}=\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})).$
@@
Then $(Z_{y},s_{y})$ is a normal crossing divisor in $(\mathbb{C}^{2},s_{y})$. In particular, $I_{s_{y}}(X_{y},W_{y})=1$.
@@
Let $M\subset\mathcal{O}_{X}^{p}$ be a finitely generated $\mathcal{O}_{X}$-module which is locally free of rank $e$ at the generic point of each irreducible component of $X$. For $A\in\text{Mat}(e\times p,\mathbb{C})$ set $[M_{e}]:=A[M]$, where $[M]$ is a presentation matrix of $\mathcal{O}_{X}^{p}/M$. Let $M_{e}\subset\mathcal{O}_{X}^{e}$ be the $\mathcal{O}_{X}$-module generated by the columns of $[M_{e}]$. Note that by the Cauchy-Binet formula $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})\subset\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$. Let $S^{\prime}$ be the analytic subspace of $X$ defined by $(\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e}):_{\mathcal{O}_{X}}\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M))$.
@@
Suppose $A$ is generic. Then $S^{\prime}$ is generically reduced.
@@
Let $\eta$ be the generic point of an irreducible component of $X$. Because $X$ is reduced, $\mathcal{O}_{X,\eta}$ is a field and so $M_{\eta}:=M\otimes\mathcal{O}_{X,\eta}$ and $(M_{e})_{\eta}:=M_{e}\otimes\mathcal{O}_{X,\eta}$ are $\mathcal{O}_{X,\eta}$-vector spaces. So $\mathcal{O}_{X,\eta}^{p}/M_{\eta}$ is $\mathcal{O}_{X,\eta}$-vector space of dimension $p-e$ and $\mathcal{O}_{X,\eta}^{e}=(M_{e})_{\eta}$. Thus the closed subsets of $X$ defined by $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$ and $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})$ are nowhere dense in $X$. Therefore, $S^{\prime}$ is either empty or of dimension $1$.
Suppose $[M]\in\text{Mat}(p\times l,\mathcal{O}_{X})$. Denote by $M^{*}$ be the $\mathcal{O}_{X}$-submodule of $\mathcal{O}_{X}^{l}$ generated by the columns of $[M]^{\mathrm{tr}}$. We have $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*})$. Denote by $M^{\prime}$ the $\mathcal{O}_{X}$-submodule of $M^{*}$ generated by the columns of $[M^{*}]A^{\text{tr}}$. Because $[M^{*}]A^{\text{tr}}=(A[M])^{\text{tr}}$, we have $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime})$.
Let $U:=X\setminus\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Then the sheaf induced by $M^{*}$ on $U$ is locally free of rank $e$. Because $X$ is reduced so is $U$. If $A$ is generic, by Kleiman’s transversality theorem [Kl74, Rmks. 6 and 7], the analytic subspace in $U$ defined by the ideal sheaf $\mathrm{Fitt}_{0}(M^{*}/M^{\prime})$ is either empty or reduced and of dimension $1$.
Suppose $S^{\prime}$ is nonempty. Let $x\neq 0$ be a point in $S^{\prime}$. Then $x\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime}))$ and $x\not\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Consider the following short exact sequence
@@
because $\mathcal{O}_{X,x}^{l}/M^{*}_{x}$ is free of rank $l-e$ so the only nonzero term in the sum above is the one corresponding to $(i,j)=(0,l-e)$. Thus locally at $x$, the subspace $S^{\prime}$ is defined in $\mathcal{O}_{X,x}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X,x}^{e}/(M_{e})_{x})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X,x}^{l}/M^{\prime}_{x})=\mathrm{Fitt}_{0}(M^{*}_{x}/M^{\prime}_{x})$. So $S^{\prime}$ is reduced at $x$. ∎
@@
If $S^{\prime}_{y}$ is empty for $y\neq 0$, then so is $S_{y}$. Thus, by Proposition 3.2 $S$ is empty. So each $X_{y}$ is therefore a complete intersection and $I(X_{y},W_{y})=0$ for all $y$. If $S^{\prime}_{y}$ is nonempty, then so is $S_{y}$. By Prp. 4.2 applied with $M=J_{\mathrm{rel}}(X)$ and $M_{e}=J_{\mathrm{rel}}(Z)$, we get that $S^{\prime}$ is generically reduced. Thus by generic smoothness $S^{\prime}_{y}$ is reduced for all $y\neq 0$ after possibly replacing $Y$ by a smaller neighborhood of $0\in Y$. By Prp. 4.1 $I_{s_{y}}(X_{y},W_{y})=1$ for each $s_{y}\in S_{y}$. Thus $S_{y}$ is reduced, too. So, $S^{\prime}_{y}=S_{y}$ for $y\neq 0$. By Proposition 3.2 and Proposition 4.1 for $y\neq 0$ we have
(7) $I_{0}(X_{0},W_{0})=I(X_{y},W_{y})=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})=\#(S_{y}).$
By Proposition 2.2 $C_{\mathrm{rel}}(X)=\mathrm{Projan}(\mathcal{R}(J_{\mathrm{rel}}(Z))$. Let $B\in\mathrm{Mat}(n\times e,\mathbb{C})$. Consider the $\mathcal{O}_{X}$-module $J_{\mathrm{rel}}(Z)^{\prime}$ generated by the columns of $[J_{\mathrm{rel}}(Z)]B$. Denote the images of $J_{\mathrm{rel}}(Z)^{\prime}$ and $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{0}}^{n-1}$ by $J(Z_{0})^{\prime}$ and $J(Z_{0})$, respectively. Choose $B$ generic so that $J(Z_{0})^{\prime}$ is a reduction $J(Z_{0})$ in $\mathcal{O}_{X_{0}}^{n-1}$. Because $J_{\mathrm{rel}}(Z)^{\prime}$ is generated by $n-1$ elements, each of which defines a generic hyperplane ˇ$\check{\mathbb{P}}^{n-1}$, then the ideal $J_{\mathrm{rel}}(Z)^{\prime}\mathcal{R}(J_{\mathrm{rel}}(Z))$ cuts out $\Gamma_{\mathrm{rel}}^{1}(X)$ before the projection to $X$. Thus $\mathrm{Supp}_{X}(J_{\mathrm{rel}}(Z)/J_{\mathrm{rel}}(Z)^{\prime})=\Gamma_{\mathrm{rel}}^{1}(X)$. The genericity condition on $B$ and Kleiman’s transversality theorem ensure that $\Gamma_{\mathrm{rel}}^{1}(X)$ meets $S$ only at 00 after possibly replacing the representative for $(X,0)$ by a smaller one.
Choosing $B$ generic enough by [Kl74, Rmks. 6 and 7] it follows that the principle ideal $(g):=\mathrm{Fitt}_{0}(\mathcal{O}_{X,0}^{n-1}/J_{\mathrm{rel}}(Z)^{\prime})$ defines a reduced subvariety of $X$, which is therefore Cohen-Macaulay, and thus it is flat over $Y$ because $Y$ is smooth. For $y\in Y$ denote the image of $g$ in $\mathcal{O}_{X_{y}}$ by $g_{y}$. By flatness $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/(g_{y}).$ But $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})$. By Proposition 2.3 we have $e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})=e(\mathrm{Jac}(X_{0}))$. Thus
@@
It remains to interpret the right-hand side of (8). We need to count the points cut out from $X_{y}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})^{\prime})$ for $y\neq 0$. There are two disjoint set of points: those that constitute the support of $J_{\mathrm{rel}}(Z_{y})/J_{\mathrm{rel}}(Z_{y})^{\prime}$ and those that are not in it. The first set of points is precisely $\Gamma_{\mathrm{rel}}^{1}(X)_{y}$. By definition the latter set is $S^{\prime}_{y}$ which is the same as $S_{y}$ as determined above. Thus
@@
Remark. A consequence of the formula we just proved is that the complete intersection discrepancy $I_{0}(X_{0},W_{0})$ of $X_{0}$ is an intrinsic invariant for a reduced smoothable $(X_{0},0)$.
@@
By Theorem 1.1 we have $e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})=\mu(X_{y})+m(X_{y})-1$, where $\mu(X_{y})$ and $m(X_{y})$ are the Milnor number and the multiplicity of $X_{y}$ at $0\in\mathbb{C}^{n}$. Because $\mu(X_{y})$ and $m(X_{y})$ are upper semi-continuous invariants the function $y\rightarrow e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})$ is constant along $Y$ if and only if $\mu(X_{y})$ and $m(X_{y})$ are constant along $Y$. Note that if $m(X_{y})$ is constant along $Y$, the singular locus does not split, i.e. the union of singularities of the fibers $X_{y}$ is $Y$. The rest of the proof follows from [BGG80, Thm. III.3].

### v3

@@ 3. A flatness result
Successive applications of prime avoidance imply that $(Z_{0},0)$ is a complete intersection for generic $A$. Next, observe that the following relation between the Jacobian matrices of $(X_{0},0)$ and $(Z_{0},0)$ holds: $J(Z_{0})_{|X_{0}}=AJ(X_{0})$. Let $(X,0)=\bigcup_{i=1}^{s}X_{i}$ be the decomposition of $(X,0)$ into irreducible components. For each $i=1,\ldots,s$ pick a smooth point $x_{i}\in X_{i}$. By the Cauchy–Binet formula for generic $A$ we have that $J(Z_{0})$ evaluated at $x_{i}$ is of maximal rank $n-1$. But $(Z_{0},0)$ is a complete intersection, so it does not have embedded components. Thus $(Z_{0},0)$ is reduced along $(X_{0},0)$. ∎
Remarks. In [BLR24, Sct. 3] we show that Bertini’s theorem yields a stronger result: for generic $A$ the complete intersection $Z_{0}$ is reduced. In [BLR24, Sct. 4] we prove that $y\rightarrow I(X_{y},W_{y})$ is constant for projective curves using a formula for the arithmetic genus of a projective curve.
@@ 4. Proofs
Let $h\colon(X,0)\rightarrow(Y,0)$ be a one-parameter embedded flat deformation of a reduced curve $(X_{0},0)\subset(\mathbb{C}^{n},0)$ with $X_{y}$ smooth for small enough $y\neq 0$. Suppose $(X,0)$ is cut out from $(\mathbb{C}^{n}\times Y,0)$ by the vanishing of the analytic functions $f_{1},\ldots,f_{p}$. Let $Z$ be the complete intersection curve in $(\mathbb{C}^{n}\times Y,0)$ with equations given by $A(f_{1},\ldots,f_{p})^{T}$ for some generic $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$. Set $W:=\overline{Z\setminus X}$ and set $S:=X\times_{Z}W$. By Prp. 3.2, $Z_{0}$ is reduced along $X_{0}$, and so is $Z$ along $X$ as remarked above. We will show below that for generic $A$ the fibers $S_{y}$ are reduced for all $y\neq 0$ close enough to 00.
Denote by $J_{\mathrm{rel}}(X)$ the relative Jacobian module and denote by $J_{\mathrm{rel}}(Z)$ the $\mathcal{O}_{X}$-module generated by the columns of the relative Jacobian matrix of $Z$ restricted to $X$. In terms of presentation matrices we have $[J_{\mathrm{rel}}(Z)]=A[J_{\mathrm{rel}}(X)]$. For each $y$ denote by $J(Z_{y})$ the image of $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{y}}^{n-1}$. It’s the $\mathcal{O}_{X_{y}}$-module generated by the columns of the Jacobian matrix of $Z_{y}$ restricted to $X_{y}$. We will repeatedly use the fact that the formation of Fitting ideals commutes with base change. In particular, $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))\otimes\mathcal{O}_{X_{y}}=\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})).$
@@
Then $(Z_{y},s_{y})$ is isomorphic to a normal crossing divisor in $(\mathbb{C}^{2},s_{y})$. In particular, we have $I_{s_{y}}(X_{y},W_{y})=1$.
@@
Let $M\subset\mathcal{O}_{X}^{p}$ be a finitely generated $\mathcal{O}_{X}$-module which is locally free of rank $e$ at the generic point of each irreducible component of $X$. For $A\in\text{Mat}(e\times p,\mathbb{C})$ set $[M_{e}]:=A[M]$, where $[M]$ is a presentation matrix of $\mathcal{O}_{X}^{p}/M$. Let $M_{e}\subset\mathcal{O}_{X}^{e}$ be the $\mathcal{O}_{X}$-module generated by the columns of $[M_{e}]$. Note that by the Cauchy-Binet formula $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})\subset\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$. Let $S^{\prime}$ be the analytic subspace of $X$ defined by $\bigcup_{i=0}^{\infty}(\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e}):_{\mathcal{O}_{X}}\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)^{i})$.
@@
Suppose $A$ is generic. Then $(S^{\prime},0)$ is reduced.
@@
Let $\eta$ be the generic point of an irreducible component of $X$. Because $X$ is reduced, $\mathcal{O}_{X,\eta}$ is a field and so $M_{\eta}:=M\otimes\mathcal{O}_{X,\eta}$ and $(M_{e})_{\eta}:=M_{e}\otimes\mathcal{O}_{X,\eta}$ are $\mathcal{O}_{X,\eta}$-vector spaces. So $\mathcal{O}_{X,\eta}^{p}/M_{\eta}$ is $\mathcal{O}_{X,\eta}$-vector space of dimension $p-e$ and $\mathcal{O}_{X,\eta}^{e}=(M_{e})_{\eta}$. Thus the closed subsets of $X$ defined by $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$ and $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})$ are nowhere dense in $X$. Therefore, $S^{\prime}$ is either empty or of dimension $1$. If $S^{\prime}$ is empty, there is nothing to prove. If $S^{\prime}$ is of dimension one, then 00 is not an embedded point of $S^{\prime}$ (see [E95, Prp. 3.13a]).
Suppose $[M]\in\text{Mat}(p\times l,\mathcal{O}_{X})$. Denote by $M^{*}$ the $\mathcal{O}_{X}$-submodule of $\mathcal{O}_{X}^{l}$ generated by the columns of $[M]^{\mathrm{tr}}$. We have $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*})$. Denote by $M^{\prime}$ the $\mathcal{O}_{X}$-submodule of $M^{*}$ generated by the columns of $[M^{*}]A^{\text{tr}}$. Because $[M^{*}]A^{\text{tr}}=(A[M])^{\text{tr}}$, we have $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime})$.
Let $x\neq 0$ be a point in $S^{\prime}$. Then $x\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime}))$ and $x\not\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Consider the following short exact sequence
@@
because $\mathcal{O}_{X,x}^{l}/M^{*}_{x}$ is free of rank $l-e$ so the only nonzero term in the sum above is the one corresponding to $(i,j)=(0,l-e)$. Thus locally at $x$, the subspace $S^{\prime}$ is defined in $\mathcal{O}_{X,x}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X,x}^{e}/(M_{e})_{x})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X,x}^{l}/M^{\prime}_{x})=\mathrm{Fitt}_{0}(M^{*}_{x}/M^{\prime}_{x})$. Let $U:=X\setminus\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Then the sheaf induced by $M^{*}$ on $U$ is locally free of rank $e$. Because $X$ is reduced so is $U$. If $A$ is generic, by Kleiman’s transversality theorem [Kl74, Rmks. 6 and 7], the analytic subspace in $U$ defined by the ideal sheaf $\mathrm{Fitt}_{0}(M^{*}/M^{\prime})$ is either empty or reduced and of dimension $1$. So $S^{\prime}$ is reduced at $x$. Thus $S^{\prime}$ satisfies Serre’s $(S_{1})$ and $(R_{0})$ conditions. So $(S^{\prime},0)$ is reduced. ∎
Let $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$. Set $M:=J_{\mathrm{rel}}(X)$ and $M_{e}:=J_{\mathrm{rel}}(Z)$. Then $S^{\prime}$ is the subscheme of $X$ cut out by $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))$. Recall that $S=X\times_{Z}W$ and by Prp. 3.1 $W_{y}=\overline{Z_{y}\setminus X_{y}}$ for all $y$. The next result shows that when $(X_{0},0)$ is smoothable we can compute the complete intersection discrepancy $I_{0}(X_{0},W_{0})$ as the number of singular points of $Z_{y}$ that are in $X_{y}$ for $y\neq 0$ generic.
Proposition 4.3.
Suppose $A$ is generic. Then $S_{y}=S^{\prime}_{y}$. In particular, $I(X_{0},W_{0})=\#(S^{\prime}_{y})$ for $y\neq 0$ close enough to 00.
Proof.
If $S^{\prime}_{y}$ is empty for $y\neq 0$, then so is $S_{y}$. Thus, by Prp. 3.1 $S$ is empty. So each $X_{y}$ is therefore a complete intersection and $I(X_{y},W_{y})=0$ for all $y$. If $S^{\prime}_{y}$ is nonempty, then so is $S_{y}$. By Prp. 4.2 applied with $M=J_{\mathrm{rel}}(X)$ and $M_{e}=J_{\mathrm{rel}}(Z)$, we get that $S^{\prime}$ is generically reduced. Thus by generic smoothness $S^{\prime}_{y}$ is reduced for all $y\neq 0$ after possibly replacing $Y$ by a smaller neighborhood of $0\in Y$. By Prp. 4.1 $I_{s_{y}}(X_{y},W_{y})=1$ for each $s_{y}\in S_{y}$. Thus $S_{y}$ is reduced, too. So, $S^{\prime}_{y}=S_{y}$ for $y\neq 0$. By Prp. 3.1 and Prp. 4.1 for $y\neq 0$ we have
(7) $I_{0}(X_{0},W_{0})=I(X_{y},W_{y})=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})=\#(S_{y})$
which completes the proof of the proposition. ∎
@@
By Prp. 2.2 $C_{\mathrm{rel}}(X)=\mathrm{Projan}(\mathcal{R}(J_{\mathrm{rel}}(Z))$. Let $B\in\mathrm{Mat}(n\times e,\mathbb{C})$. Consider the $\mathcal{O}_{X}$-module $J_{\mathrm{rel}}(Z)^{\prime}$ generated by the columns of $[J_{\mathrm{rel}}(Z)]B$. Denote the images of $J_{\mathrm{rel}}(Z)^{\prime}$ and $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{0}}^{n-1}$ by $J(Z_{0})^{\prime}$ and $J(Z_{0})$, respectively. Choose $B$ generic so that $J(Z_{0})^{\prime}$ is a reduction $J(Z_{0})$ in $\mathcal{O}_{X_{0}}^{n-1}$. Because $J_{\mathrm{rel}}(Z)^{\prime}$ is generated by $n-1$ elements, each of which defines a generic hyperplane ˇ$\check{\mathbb{P}}^{n-1}$, then the ideal $J_{\mathrm{rel}}(Z)^{\prime}\mathcal{R}(J_{\mathrm{rel}}(Z))$ cuts out $\Gamma_{\mathrm{rel}}^{1}(X)$ before the projection to $X$. Thus $\mathrm{Supp}_{X}(J_{\mathrm{rel}}(Z)/J_{\mathrm{rel}}(Z)^{\prime})=\Gamma_{\mathrm{rel}}^{1}(X)$. The genericity condition on $B$ and Kleiman’s transversality theorem ensure that $\Gamma_{\mathrm{rel}}^{1}(X)$ meets $S$ only at 00 after possibly replacing the representative for $(X,0)$ by a smaller one.
Choosing $B$ generic enough by [Kl74, Rmks. 6 and 7] it follows that the principle ideal $(g):=\mathrm{Fitt}_{0}(\mathcal{O}_{X,0}^{n-1}/J_{\mathrm{rel}}(Z)^{\prime})$ defines a reduced subvariety of $X$, which is therefore Cohen-Macaulay, and thus it is flat over $Y$ because $Y$ is smooth. For $y\in Y$ denote the image of $g$ in $\mathcal{O}_{X_{y}}$ by $g_{y}$. By flatness $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/(g_{y}).$ But $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})$. By Prp. 2.3 we have $e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})=e(\mathrm{Jac}(X_{0}))$. Thus
@@
It remains to interpret the right-hand side of (8). We need to count the points cut out from $X_{y}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})^{\prime})$ for $y\neq 0$. There are two disjoint set of points: those that constitute the support of $J_{\mathrm{rel}}(Z_{y})/J_{\mathrm{rel}}(Z_{y})^{\prime}$ and those that are not in it. The first set of points is precisely $\Gamma_{\mathrm{rel}}^{1}(X)_{y}$. By definition the latter set is $S^{\prime}_{y}$ which is the same as $S_{y}$ as shown in Prp. 4.3. Thus
@@
Remarks. By Prp. 3.1 $W_{0}=\overline{Z_{0}\setminus X_{0}}$. Thus a consequence of the formula we just proved is that the complete intersection discrepancy $I_{0}(X_{0},W_{0})$ of $X_{0}$ is an intrinsic invariant for a reduced smoothable $(X_{0},0)$ because $\mu$, $m$ and $e(\mathrm{Jac}(X_{0}))$ are intrinsic invariants of $(X_{0},0)$.
Let $F\colon(\mathbb{C}^{n},0)\rightarrow(\mathbb{C},0)$ be a general linear functional. In [NT08, Cor. 3.2] it’s shown that $\mu(F|X_{0})=\mu+m-1$ where $\mu(F|X_{0})$ is the Milnor number of $F|X_{0}$. Thus by Thm. 1.1 we obtain that $\mu(F|X_{0})=e(\mathrm{Jac}(X_{0}))-I_{0}(X_{0},W_{0})$. This formula can be used to compute $\mu(F|X_{0})$ in case the parametrization of $X_{0}$ is not known.
@@
By Theorem 1.1 we have $e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})=\mu(X_{y})+m(X_{y})-1$, where $\mu(X_{y})$ and $m(X_{y})$ are the Milnor number and the multiplicity of $X_{y}$ at $0\in\mathbb{C}^{n}$. Because $\mu(X_{y})$ and $m(X_{y})$ are upper semi-continuous invariants (see [BG80, Thm. 6.1.7.(b)] for the upper semi-continuity of $\mu$) the function $y\rightarrow e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})$ is constant along $Y$ if and only if $\mu(X_{y})$ and $m(X_{y})$ are constant along $Y$. Note that if $\mu(X_{y})$ is constant along $Y$, by [BG80, Lem. 5.2.1] the singular locus does not split, i.e. the union of singularities of the fibers $X_{y}$ is $Y$. The rest of the proof follows from [BGG80, Thm. III.3] (see also [NT08, Thm. 4.3]).